# Business Informatics Research Center, Cluj-Napoca, Romania

### v1

@@ 4 Results and discussion
Implicit reasoning poses challenges for LLMs. Template Hard (TH) dataset contains cases where the LLM needs to understand that a given value is already an ID that references an existing instance in a KG or that a statement implies a specific relationship pertaining to a class. As concluded by the results preented in the tables, under flexible metrics, Mixtral 8x7B achieves an accuracy of and an F1 score of on the more difficult dataset, which is lower than its performance on the easier one. GPT-3.5 Turbo narrows this margin, reducing from a peak accuracy and F1 score of to on Template Hard (TH). Interestingly enough, Mixtral 8x7B yields its best scores at level 1 prompts, when strictly measured.
@@
In summary, KGC remains a challenging task for LLMs under Zero-Shot prompting. Moreover, when checking their intermediate reasoning steps, they show lack of ability to follow the provided ontology. The open-source model has difficulties in conforming to the required output format. However, One-Shot contexts give promising results as LLMs excel in emulating a provided example. This implies that a less resource-intensive Few-Shot training approach could potentially boost performances, with a focus on techniques like Retrieval-Augmented-Generation to select more suitable examples within a given prompt. Another plus is their ability to enhance their inner knowledge to detect some implicit relationships from the input text. Nevertheless, as suggested by Fill et al. [1], presently we may use such LLMs as helpful assistants for solving such tasks, rather than ultimately faithful extractors in a pipelined system.
@@ 5 Conclusion
The proposed experiments showcases the ability of two leading LLMs, namely Mixtral-8x7B-Instruct-v0.1 and gpt-3.5-turbo-0125, in tackling the Knowledge Graph Completion task. Using both hand-written and model-rephrased prompts, we incorporated various prompt engineering techniques, such as In-Context Learning or Chain of Thought, focusing on Zero- and One-Shot contexts. Metrics measurement enabled the evaluation of the LLM strictly following the given prompt, as well as the its flexibility in considering post-processing steps. The results obtained from two distinct datasets tailored to various reasoning challenges highlight the LLMs strengths and weaknesses. These include their adaptability in Zero- or One-Shot scenarios and their utilization of internal knowledge to deduce implicit reasoning steps. However, they still lack self-awareness, not being able to adhere to explicit guidelines in the given prompt, or fully understand and exploit the considered ontology.

### v2

@@ 4 Results and discussion
Implicit reasoning poses challenges for LLMs. Template Hard (TH) dataset contains cases where the LLM needs to understand that a given value is already an ID that references an existing instance in a KG or that a statement implies a specific relationship pertaining to a class. As concluded by the results presented in the tables, under flexible metrics, Mixtral 8x7b achieves an accuracy of and an F1 score of on the more difficult dataset, which is lower than its performance on the easier one. GPT-3.5-Turbo narrows this margin, reducing from a peak accuracy and F1 score of to on Template Hard (TH). Same behaviour is observed with GPT-4o, as it falls from accuracy and F1 score to around . Interestingly enough, Mixtral-8x7b yields its best scores at level 1 prompts, when strictly measured.
GPT-4o is more consistent and performant, while GPT-3.5-Turbo achieves the best results on the harder dataset. Despite showing fluctuations in results when it rephrased the prompts, GPT-4o was the best overall model. On the TE dataset, under strict measurements, it had accuracy and F1 Score, almost four times more than Mixtral-8x7b and with more than GPT-3.5-Turbo. We can interpret the results as GPT-4o is more reliable than the other two models, regardless of the prompt level. However, GPT-3.5-Turbo came close to it considering their top performances, being only away from GPT-4o on the TE dataset, while surpassing it by on the TH dataset. Depending on the user’s objectives, while considering the model’s costs, the choice of the final model could vary.
@@
In summary, KGC remains a challenging task for LLMs under Zero-Shot prompting. As models become better, they performance tend to increase, while shifting the focus on optimizing the costs. Moreover, when checking their intermediate reasoning steps, they show lack of ability to follow the provided ontology. The open-source model has difficulties in conforming to the required output format. However, One-Shot contexts give promising results as LLMs excel in emulating a provided example. This implies that a less resource-intensive Few-Shot training approach could potentially boost performances, with a focus on techniques like Retrieval-Augmented-Generation to select more suitable examples within a given prompt. Another plus is their ability to enhance their inner knowledge to detect some implicit relationships from the input text. Nevertheless, as suggested by Fill et al. [2], presently we may use such LLMs as helpful assistants for solving such tasks, rather than ultimately faithful extractors in a pipelined system.
@@ 5 Conclusion
The proposed experiments showcases the ability of three leading LLMs, namely Mixtral-8x7b-Instruct-v0.1, GPT-3.5-Turbo-0125 and GPT-4o, in tackling the Knowledge Graph Completion task. Using both hand-written and model-rephrased prompts, we incorporated various prompt engineering techniques, such as In-Context Learning or Chain of Thought, focusing on Zero- and One-Shot contexts. Metrics measurement enabled the evaluation of the LLM strictly following the given prompt, as well as the its flexibility in considering post-processing steps. The results obtained from two distinct datasets tailored to various reasoning challenges highlight the LLMs strengths and weaknesses. These include their adaptability in Zero- or One-Shot scenarios and their utilization of internal knowledge to deduce implicit reasoning steps. However, they still lack self-awareness, not being able to adhere to explicit guidelines in the given prompt, or fully understand and exploit the considered ontology.